# A Fine-Grained Attention and Geometric Correspondence Model for Musculoskeletal Risk Classification in Athletes Using Multimodal Visual and Skeletal Features

### v1

@@ 3.2  Proposed model
Figure 3 illustrates the proposed ViSK-GAT framework. The most significant contribution of the framework is the Multimodal Geometric Correspondence Module (MGCM), which aligns visual features with coordinate-based skeletal representations through cross-attention mechanisms (see Section 3.2.3). MGCM maps both modalities into a shared latent space, where attention-based feature alignment allows the model to learn complex cross-modal correspondences and geometric relationships. This alignment facilitates a coherent understanding of spatial structures and visual semantics. We further enhance the integrated features using a final transformer-based refinement stage (see Section 3.2.1), which captures long-range contextual dependencies across modalities. This leads to a unified and discriminative multimodal representation suitable for complex musculoskeletal risk classification tasks.
@@ 3.2.3  Multimodal Geometric Correspondence Module
Figure 5: Proposed architecture of the geometric correspondence module
@@
where ${F}^{\prime}_{\text{img}}\in\mathbb{R}^{256\times d}$ serves as the query matrix, and ${F}^{\prime}_{\text{pose}}\in\mathbb{R}^{33\times d}$ is used as both the key and value matrices. We then concatenated ${F}^{\prime}_{\text{img}}$ with ${F}_{\text{attn}}$ to form a fused representation ${F}_{\text{fused}}\in\mathbb{R}^{1\times 2d}$. To capture higher-order cross-modal interactions and enrich geometric understanding, we processed ${F}_{\text{fused}}$ through 2 stacked Transformer layers. Each layer consisted of multi-head self-attention with 4 heads, feed-forward blocks with dimensionality expanded to $4d$, and residual normalization. This transformer stack refined the fused representation and modeled long-range dependencies across modalities. Finally, the output token was passed through a fusion head comprising a linear layer, LayerNorm, and GELU activation to generate the joint embedding ${F}_{\text{corr}}\in\mathbb{R}^{512}$, used for final classification. This operation is summarized as shown in Equation (7):
@@ 4.2  Training analysis
The training process was carried out using a supervised approach on the proposed multimodal dataset. The full dataset contained image-coordinate pairs annotated with REBA-based posture risk labels. The dataset was divided into 70% training, 10% validation, and 20% testing sets. We trained our multimodal model for 100 epochs using a batch size of 16. The optimizer used was AdamW with a weight decay of $1\times 10^{-5}$ to prevent overfitting. To improve learning dynamics, we used the OneCycleLR learning rate scheduler. This scheduler started with a low learning rate, gradually increased to a peak, and then decreased again over the course of training. The peak learning rate was set to $3\times 10^{-4}$, with a warm-up over 10% of the epochs and a final decay according to a div factor of 1000. The model was trained on the resized ($224\times 224$) image data using a categorical cross-entropy loss function. At the end of training, the model achieved a training loss of 0.0007 and a training accuracy of 99.97% in the final epoch.
@@ 4.3  Classification assessment
Class-wise Performance. The model demonstrates consistently strong classification ability across all REBA classes, with F1-scores ranging from 88% to 99%. Table 2 summarizes these per-class performance metrics.
Table 2: Class-wise precision, recall, F1-score, and specificity for REBA classification
@@
In contrast, classes 2 and 4 exhibit comparatively lower F1-scores of 89.32% and 88.58%, respectively, as these moderate-risk classes contain more ambiguous or transitional postures, which increase confusion with adjacent categories. We note that precision for class 2 is the lowest overall (88.46%), and recall for class 4 is the lowest (87.70%), suggesting these classes are the most challenging for the model. We also analyzed the multiclass ROC curves to further investigate the model’s discriminative ability across the eight REBA classes.
As shown in Figure 6 (left), the model shows excellent class separability, with most classes achieving an Area Under the Curve (AUC) of 1.0. Classes 2 and 3 scored slightly lower AUCs of 0.99, while class 4 recorded an AUC of 0.98. This observation is further supported by the confusion matrix in Figure 6 (right).
Overall Evaluation. The model demonstrates superior musculoskeletal risk classification with an accuracy approaching 94%. However, because accuracy alone can be misleading in multiclass settings, especially if class distribution is imbalanced, we further consider Cohen’s Kappa and the Matthews Correlation Coefficient (MCC), both of which are 93%. These metrics adjust for chance agreement and class imbalance, respectively. Additionally, the predicted probability distribution’s RMSE and the MAE are 0.1205 and 0.0156, respectively. These low values indicate that the model’s predicted probability distributions are highly confident and closely aligned with the true labels, even for misclassified samples. The overall metrics are presented in Table 3.
Table 3: Overall classification metrics of the proposed model.
@@
Secondary Error Analysis. To identify where misclassifications are concentrated in the experiment, we examined Negative Predictive Value (NPV), False Positive Rate (FPR), False Discovery Rate (FDR), and False Negative Rate (FNR) for each class.
@@
Class 1 0.99.75 0.0057 0.0415 0.0189
Class 2 0.98.38 0.0194 0.1154 0.0980
Class 3 0.98.72 0.0135 0.0875 0.0837
Class 4 0.97.36 0.0169 0.1053 0.1565
Class 5 0.99.68 0.0070 0.0451 0.0210
Class 6 0.99.24 0.0057 0.0393 0.0517
Class 7 0.99.88 0.0012 0.0121 0.0121
Class 8 0.99.94 0.0012 0.0102 0.0051
@@
[…]
Baseline Simple Concat 0.8712 0.867 N/A
+ ResBlocks Simple Concat 0.9028 0.891 +0.0316
+ FGAM Simple Concat 0.9213 0.911 +0.0501
+ MGCM (ViSK-GAT) Attention-Fuse 0.9385 0.930 +0.0673
Table 6: Ablation study on training hyperparameters: learning rate (LR), batch size, weight decay, optimizer, gradient clipping threshold (Clip), label smoothing (L/ Smooth), and loss function (Loss Func). Training time is reported in minutes per epoch (min/Eps). Best values in the training accuracy and loss are presented in bold, and the selected configuration for our final training is highlighted.
Exp LR Batch Weight Decay Optimizer Clip L/ Smooth Loss Func min/Ep Accuracy Loss
1 1e-3 32 1e-3 Adam 1.0 0.1 CrossEntropy $\sim$2:13 0.0018 0.9623
2 1e-4 16 1e-4 Adam 1.0 0.1 Focal $\sim$1:45 0.0042 0.9380
3 2e-4 32 1e-5 SGD 0.5 0.1 Focal $\sim$0:50 0.0060 0.9462
4 4e-4 16 1e-5 Adam 1.0 0.2 CE + Focal $\sim$1:57 0.0014 0.9665
5 3e-4 16 1e-5 AdamW 1.0 0.1 CrossEntropy $\sim$1:34 0.0007 0.9997
6 3e-4 32 1e-4 Adam 2.0 0.1 CrossEntropy $\sim$0:51 0.0017 0.9584
7 4e-4 16 1e-4 AdamW 1.0 0.1 Focal $\sim$1:50 0.0010 0.9822
8 3e-4 8 1e-5 Adam 1.0 0.1 CrossEntropy $\sim$2:40 0.0021 0.9995
9 4e-4 32 1e-5 AdamW 1.0 0.1 CE + Focal $\sim$1:10 0.0018 0.9745
10 3e-4 32 1e-5 AdamW 0.5 0.2 CrossEntropy $\sim$0:58 0.0012 0.9691
11 4e-4 16 1e-4 Adam 1.0 0.1 CrossEntropy $\sim$1:35 0.0015 0.9630
12 3e-4 8 1e-5 SGD 1.0 0.1 CE + Focal $\sim$2:20 0.0050 0.9342
4.4  Ablation studies
[…]
Table 7: Comparison of our proposed multimodal model with various standard image feature extractor backbones combined with an MLP for coordinate data. All models use a common fusion and classification pipeline.
Backbone Val Acc (%) Val Loss Test Acc (%) Test Loss
VGG16 83.04 0.6985 83.13 0.7273
VGG19 84.12 0.6844 84.20 0.6911
ResNet50 87.81 0.5981 87.75 0.6746
ResNet101 88.44 0.5832 88.39 0.5812
InceptionV3 87.03 0.6114 87.02 0.6105
Xception 88.06 0.5908 88.13 0.5909
EfficientNet-B0 87.45 0.6043 87.42 0.6129
MobileNetV2 84.25 0.6732 84.23 0.6714
DenseNet121 88.12 0.5925 88.06 0.5947
ViSK-GAT 93.66 0.4376 93.89 0.4808
Table 8: Comparison with prior studies in ergonomic posture risk assessment.
@@ 4.4  Ablation studies
Zhao et al. [25] Multimodal Collected 95.0% N/A
@@
ViSK-GAT Multimodal Constructed 93.89% REBA
Hyperparameter Configurations. We conducted an extensive ablation study of 12 different experiments (Exp.) to evaluate the impact of various training hyperparameters on our multimodal posture classification model. The parameters tested include learning rate, batch size, weight decay, optimizer type, gradient clipping threshold, label smoothing, and loss functions.
In terms of loss functions, standard CrossEntropy consistently performed well, whereas the combination with Focal loss (Exps. 4, 9, 12) slightly enhanced robustness in some cases but incurred longer training times. Using Focal loss alone (Exps. 2, 3, 7) helped with class imbalance to some extent, but was sensitive to the choice of optimizer. Moderate label smoothing (0.1) proved sufficient across most runs, while higher smoothing (0.2 in Exps. 4, 10) only offered marginal benefit. Similarly, clipping gradients above 1.0 (Exp. 6) led to reduced regularization and poorer generalization. Exp. 5, which employed a learning rate of 3e-4, AdamW, batch size 16, and CrossEntropy loss with 0.1 smoothing and clipping of 1.0, offered the best trade-off and achieved high training accuracy among all of the tested settings. Table 6 summarizes these configurations, along with corresponding training times per epoch, training and validation losses, and accuracies.
@@ 4.5  Comparison with state-of-the-art models
4.6  Comparison with existing studies
@@ 5  Discussion
We used the REBA scoring framework in our classification task to ensure ergonomic relevance and interpretability. We extracted 33 anatomical key points and 2D skeletal coordinates using MediaPipe. From these keypoints, we computed the joint angles using cosine and slope-based methods. These angles were then mapped to region-specific REBA posture scores, which were combined using standard REBA lookup tables [45] to compute a final risk score. Based on these scores, each sample was classified into one of eight discrete REBA risk classes, with higher class numbers indicating higher ergonomic risk.

### v2

@@ 3.2  Proposed model
The most significant contribution of the framework is the Multimodal Geometric Correspondence Module (MGCM), which aligns visual features with coordinate-based skeletal representations through cross-attention mechanisms (see Section 3.2.3). MGCM maps both modalities into a shared latent space, where attention-based feature alignment allows the model to learn complex cross-modal correspondences and geometric relationships. This alignment facilitates a coherent understanding of spatial structures and visual semantics. We further enhance the integrated features using a final transformer-based refinement stage (see Section 3.2.1), which captures long-range contextual dependencies across modalities. This leads to a unified and discriminative multimodal representation suitable for complex musculoskeletal risk classification tasks. Figure 3 illustrates the proposed ViSK-GAT framework.
@@ 3.2.2  Fine-grained attention module
Figure 5: Proposed architecture of the Multimodal Geometric Correspondence Module (MGCM). This is a cross-attention transformer designed to bridge the gap between image tokens (Query) and skeletal coordinate embeddings (Key and Value). The process involves: (1) projecting both features into a common latent space; (2) using cross-attention to align visual appearance with structural posture information; (3) concatenating the attention output with the image feature; and (4) refining the combined feature using stacked Transformer layers to produce the final Fused Feature Vector ($F_{corr}$) for risk classification.
@@ 3.2.3  Multimodal geometric correspondence module
where ${F}^{\prime}_{\text{img}}\in\mathbb{R}^{256\times d}$ serves as the query matrix, and ${F}^{\prime}_{\text{pose}}\in\mathbb{R}^{33\times d}$ is used as both the key and value matrices. We then concatenated ${F}^{\prime}_{\text{img}}$ with ${F}_{\text{attn}}$ to form a fused representation ${F}_{\text{fused}}\in\mathbb{R}^{1\times 2d}$. To capture higher-order cross-modal interactions and enrich geometric understanding, we processed ${F}_{\text{fused}}$ through 2 stacked Transformer layers. Each layer consisted of multi-head self-attention with 4 heads, feed-forward blocks with dimensionality expanded to $4d$, and residual normalization.
This transformer stack refined the fused representation and modeled long-range dependencies across modalities. Finally, the output token was passed through a fusion head comprising a linear layer, LayerNorm, and GELU activation to generate the joint embedding ${F}_{\text{corr}}\in\mathbb{R}^{512}$, used for final classification. This operation is summarized as shown in Equation (7):
@@ 4.2  Training analysis
The training process was carried out using a supervised approach on the proposed multimodal dataset. The full dataset contained image-coordinate pairs annotated with REBA-based posture risk labels. The dataset was divided into 70% training, 10% validation, and 20% testing sets. We trained our multimodal model for 100 epochs using a batch size of 16. The optimizer used was AdamW with a weight decay of $1\times 10^{-5}$ to prevent overfitting. To improve learning dynamics, we used the OneCycleLR learning rate scheduler. This scheduler started with a low learning rate, gradually increased to a peak, and then decreased again over the course of training. The peak learning rate was set to $3\times 10^{-4}$, with a warm-up over 10% of the epochs and a final decay according to a div factor of 1000. The model was trained on the resized ($224\times 224$) image data using a categorical cross-entropy loss function. At the end of training, the model achieved a training loss of 0.0007 and a training accuracy of 95.97% in the final epoch.
@@ 4.3  Classification assessment
4.3.1  Class-wise performance
The model demonstrates consistently strong classification ability across all REBA classes, with F1-scores ranging from 88% to 99%. Table 2 summarizes these per-class performance metrics.
Table 2: Class-wise precision, recall, F1-score, and specificity for REBA classification.
@@ 4.3.1  Class-wise performance
In contrast, classes 2 and 4 exhibit comparatively lower F1-scores of 89.32% and 88.58%, respectively, as these moderate-risk classes contain more ambiguous or transitional postures, which increase confusion with adjacent categories. We note that precision for class 2 is the lowest overall (88.46%), and recall for class 4 is the lowest (87.70%), suggesting these classes are the most challenging for the model. We also analyzed the multiclass ROC curves to further investigate the model’s discriminative ability across the eight REBA classes. As shown in Figure 6 (left), the model shows excellent class separability, with most classes achieving an AUC of 1.0. Classes 2 and 3 scored slightly lower AUCs of 0.99, while class 4 recorded an AUC of 0.98. This observation is further supported by the confusion matrix in Figure 6 (right).
4.3.2  Overall evaluation
The model demonstrates superior musculoskeletal risk classification with an accuracy approaching 94%. However, because accuracy alone can be misleading in multiclass settings, especially if class distribution is imbalanced, we further consider Cohen’s Kappa and the Matthews Correlation Coefficient (MCC), both of which are 93%. These metrics adjust for chance agreement and class imbalance, respectively. These low values indicate that the model’s predicted probability distributions are highly confident and closely aligned with the true labels, even for misclassified samples. The overall predicted probability distribution and result metrics are presented in Table 3.
Table 3: Overall classification metrics along with predicted error probability distribution of the proposed model.
@@ 4.3.2  Overall evaluation
4.3.3  Secondary error analysis
To identify where misclassifications are concentrated in the experiment, we examined Negative Predictive Value (NPV), False Positive Rate (FPR), False Discovery Rate (FDR), and False Negative Rate (FNR) for each class.
@@ 4.3.3  Secondary error analysis
Class 1 0.9975 0.0057 0.0415 0.0189
Class 2 0.9838 0.0194 0.1154 0.0980
Class 3 0.9872 0.0135 0.0875 0.0837
Class 4 0.9736 0.0169 0.1053 0.1565
Class 5 0.9968 0.0070 0.0451 0.0210
Class 6 0.9924 0.0057 0.0393 0.0517
Class 7 0.9988 0.0012 0.0121 0.0121
Class 8 0.9994 0.0012 0.0102 0.0051
@@
4.4  Comparison with existing studies
Table 5: Comparison of the proposed ViSK-GAT framework with prior studies in ergonomic posture risk assessment. The table highlights key methodological attributes, including the use of sensors, multimodal data fusion, and the ergonomic scoring system (REBA, RULA, OWAS).
@@ 4.4  Comparison with existing studies
Zhao et al. [26] Multimodal Collected 95.00% N/A
@@
ViSK-GAT (ours) Multimodal Constructed 93.89% REBA
@@ 4.5  Comparison with state-of-the-art models
Table 6: Comparison of our proposed multimodal model with various standard image feature extractor backbones combined with an MLP for coordinate data.
Backbone Val Accuracy Val Loss Test Accuracy Test Loss
VGG16 83.04% 0.6985 83.13% 0.7273
VGG19 84.12% 0.6844 84.20% 0.6911
ResNet50 87.81% 0.5981 87.75% 0.6746
ResNet101 88.44% 0.5832 88.39% 0.5812
InceptionV3 87.03% 0.6114 87.02% 0.6105
Xception 88.06% 0.5908 88.13% 0.5909
EfficientNet-B0 87.45% 0.6043 87.42% 0.6129
MobileNetV2 84.25% 0.6732 84.23% 0.6714
DenseNet121 88.12% 0.5925 88.06% 0.5947
ViSK-GAT (ours) 93.66% 0.4376 93.89% 0.4808
@@
4.6  Ablation studies
4.6.1  Module configurations
[…]
Baseline 0.8712 0.867 N/A
+ ResBlocks 0.9028 0.891 +0.0316
+ FGAM 0.9213 0.911 +0.0501
+ MGCM (ViSK-GAT) 0.9385 0.930 +0.0673
Table 8: Ablation study on training hyperparameters: learning rate (LR), batch size, weight decay, optimizer, gradient clipping threshold (Clip), label smoothing (L/ Smooth), and loss function (Loss). Training inference time is reported in minutes per epoch (Inf. Time). Best values in the training accuracy and loss are presented in bold, and the selected configuration for our final training is highlighted.
Exp LR Batch Weight Decay Optimizer Clip L/ Smooth Loss Inf. Time Loss Accuracy
1 1e-3 32 1e-3 Adam 1.0 0.1 CrossEntropy $\sim$2:13 0.0018 0.9323
2 1e-4 16 1e-4 Adam 1.0 0.1 Focal $\sim$1:45 0.0042 0.9080
3 2e-4 32 1e-5 SGD 0.5 0.1 Focal $\sim$0:50 0.0060 0.9162
4 4e-4 16 1e-5 Adam 1.0 0.2 CE + Focal $\sim$1:57 0.0014 0.9365
5 3e-4 16 1e-5 AdamW 1.0 0.1 CrossEntropy $\sim$1:34 0.0007 0.9597
6 3e-4 32 1e-4 Adam 2.0 0.1 CrossEntropy $\sim$0:51 0.0017 0.9284
7 4e-4 16 1e-4 AdamW 1.0 0.1 Focal $\sim$1:50 0.0010 0.9522
8 3e-4 8 1e-5 Adam 1.0 0.1 CrossEntropy $\sim$2:40 0.0021 0.9995
9 4e-4 32 1e-5 AdamW 1.0 0.1 CE + Focal $\sim$1:10 0.0018 0.9945
10 3e-4 32 1e-5 AdamW 0.5 0.2 CrossEntropy $\sim$0:58 0.0012 0.9391
11 4e-4 16 1e-4 Adam 1.0 0.1 CrossEntropy $\sim$1:35 0.0015 0.9230
12 3e-4 8 1e-5 SGD 1.0 0.1 CE + Focal $\sim$2:20 0.0050 0.9042
4.6.2  Hyperparameter Configurations
We conducted an extensive ablation study of 12 different experiments (Exp.) to evaluate the impact of various training hyperparameters on our multimodal posture classification model. The parameters tested include learning rate, batch size, weight decay, optimizer type, gradient clipping threshold, label smoothing, and loss functions.
In terms of loss functions, standard CrossEntropy consistently performed well, whereas the combination with Focal loss (Exps. 4, 9, 12) slightly enhanced robustness in some cases but incurred longer training times. Using Focal loss alone (Exps. 2, 3, 7) helped with class imbalance to some extent, but was sensitive to the choice of optimizer. Moderate label smoothing (0.1) proved sufficient across most runs, while higher smoothing (0.2 in Exps. 4, 10) only offered marginal benefit. Similarly, clipping gradients above 1.0 (Exp. 6) led to reduced regularization and poorer generalization. Exp. 5, which employed a learning rate of 3e-4, AdamW, batch size 16, and CrossEntropy loss with 0.1 smoothing and clipping of 1.0, offered the best trade-off accuracy and achieved the lowest training loss among all of the settings. Table 8 summarizes these configurations, along with corresponding training times per epoch, training and validation losses, and accuracies.
4.6.3  Features fusion
To further validate our architecture, we compared ViSK-GAT against standard multimodal fusion baselines, holding the training data and hyperparameters constant. As shown in Table 8, simple fusion strategies like early and late fusion perform poorly, as they cannot model complex interactions. A standard cross-attention baseline performs better but is still outperformed by ViSK-GAT. This demonstrates that the specific, sequential refinement of FGAM followed by the geometric alignment of MGCM provides a significant advantage over a naive cross-modal approach. The detailed findings are reported in Table 9.
Table 9: Ablation of the fusion module. The results are compared with existing early fusion, late fusion, and the standard transformer encoder fusion strategies.
Strategy Fusion Method Accuracy F1-Score
Early Fusion Pixel-coordinate concatenation at input 71.31% 68.91%
Late Fusion Concatenation of final image/pose features 84.23% 83.01%
Cross-Attention Standard transformer encoder fusion 88.77% 86.51%
ViSK-GAT (ours) FGAM + MGCM 93.89% 93.85%
4.6.4  Model complexity analysis
We conducted a comprehensive analysis of our model’s architectural footprint (see experimental hardware configuration in Section 4.2) to assess its suitability for real-world applications. The complete model comprises a total of 5.8 million parameters. The model requires 22.8 MB for trainable parameters and only 231 KB for non-trainable components, which results in a compact memory footprint suitable for various hardware platforms.
Table 10: Computational profile of the proposed ViSK-GAT model in terms of the major modules, including the Fine-Grained Attention Module (FGAM), and the Multimodal Geometric Correspondence Module (MGCM).
Component Parameter Trainable Non-Trainable
Image Encoder 3.5 M 14.2 MB 121 KB
Coordinate Encoder 0.1 M 0.4 MB 8 KB
FGAM 1.2 M 4.8 MB 63 KB
MGCM 0.9 M 3.2 MB 39 KB
Classification Head 0.1 M 0.2 MB 0 KB
Total 5.8 M 22.8 MB 231 KB
The model demonstrated efficient training processing a full epoch in 1:34 minutes (see inference details in Table 8). The inference speed was measured at 8.2 milliseconds per frame, which translates to a processing rate of 122 frames per second. As detailed in Table 10, the efficient training time, compact memory footprint, and rapid inference speed demonstrate that the model is computationally feasible for both development and deployment phases in real-time monitoring and assessment.
@@ 5  Discussion
We used the REBA scoring framework in our classification task to ensure ergonomic relevance and interpretability. To create a system capable of scalable deployment in diverse athletic environments, we based our analysis on 2D skeletal coordinates extracted using MediaPipe. This approach was strategically chosen because it allows the model to operate on standard monocular video, which is vastly more accessible and practical in real-world sports settings than complex, laboratory-bound 3D motion capture systems. From the 33 extracted 2D keypoints, we computed joint angles using cosine and slope-based methods. These angles were then mapped to region-specific REBA posture scores, which were combined using standard REBA lookup tables [46] to compute a final risk score. While this 2D approximation introduces a known source of variation compared to a 3D biomechanical analysis, it provides a consistent and computationally efficient proxy for postural stress. Based on these scores, each sample was classified into one of eight discrete REBA risk classes, with higher class numbers indicating higher ergonomic risk.